OpenStudy (anonymous):

A firm had retained earnings of $100,000 in 2009 and $125,000 in 2010. The increase in retained earnings from 2009 to 2010 is: A. 12.5 percent. B. 20 percent. C. 25 percent. D. 125 percent.

OpenStudy (anonymous):

right answer is 12.5 percent

OpenStudy (jdoe0001):

it's? ok, if you say so :S

OpenStudy (anonymous):

sure this is right answer iam confuse

OpenStudy (jdoe0001):

let's see if 100,000 is the 100%, the new retained earnings is 125,000 so it's really 25,000 more than before so how much is 25,000 of 100,000? or lemme put it differently \(\begin{matrix} \$ && \%\\ \cfrac{100,000}{125,000}& =& \cfrac{100}{x} \end{matrix}\)

OpenStudy (jdoe0001):

solve for "x" to find out how much 125,000 is in % compared to 100,000

OpenStudy (jdoe0001):

or we can also use \(\begin{matrix} \$ && \%\\ \cfrac{100,000}{25,000}& =& \cfrac{100}{x} \end{matrix}\) to check how much is 25,000 in % compared to 100,000 either way

OpenStudy (anonymous):

i am do not understand please help me

OpenStudy (jdoe0001):

well, just cross-multiply to solve for "x"

OpenStudy (anonymous):

answer is 25 percent

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

i am right or no please tell me

OpenStudy (jdoe0001):

well, did you solve the fractions? what did they give you?

OpenStudy (jdoe0001):

say using this one \(\begin{matrix} \$ && \%\\ \cfrac{100,000}{25,000}& =& \cfrac{100}{x} \end{matrix}\)

OpenStudy (anonymous):

25 percent

OpenStudy (jdoe0001):

if that's what you got, then that's it

OpenStudy (anonymous):

cross multiply

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

can i ask u few question more please reply me

OpenStudy (jdoe0001):

it's easier if you post anew, so we can all see and help and revise each other :)

OpenStudy (anonymous):

my question how is measure accounts payable $ 140,000 bonds payable,2020 100,000 common stock ($25 par value) 200,000 retained earnings 230,000 how is the measure the ratio o stock holders equity to total liabilities is

OpenStudy (anonymous):

method is stockholders equity / total liabilities but i do not know about to stock holders equity and total liabilities